OpenStudy (anonymous):

Andrew keeps 1/12 of his savings under his pillow. If he has $31 under his pillow, how much has Andrew saved?

TheSmartOne (thesmartone):

Welcome to Open Study!!

OpenStudy (anonymous):

372 dollars

TheSmartOne (thesmartone):

You would need to put it in a proportion\[\frac{ 1 }{ 12 }=\frac{ 31 }{ x }\]

TheSmartOne (thesmartone):

Then you would cross multiply to get your answer.

OpenStudy (anonymous):

Ya and then the answer is 372 dumkopf!

TheSmartOne (thesmartone):

\[1 \times x= 12 \times 31\]

TheSmartOne (thesmartone):

@bmjvirtual123 Read the Code of Conduct. You have to explain how you got your answer.

OpenStudy (anonymous):

I got it before you even though u r a mathlete…… shows…..

OpenStudy (anonymous):

sorry thesmartone will do

OpenStudy (anonymous):

i am new here

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

UR WELCOME ;)

TheSmartOne (thesmartone):

No problem. I was new here once too and made the same mistake...

OpenStudy (anonymous):

Ya I just made my account yesterday lol

TheSmartOne (thesmartone):

@bmjvirtual123 \[\href{http:///openstudy.com/code-of-conduct}{\large \sf Code~of~Conduct}\]